OpenStudy (anonymous):

exact value if a=0 and b =1 evaluate the integral: (9x^3 . (sqroot x) -2x^5 + e^-2x +11x) dx

OpenStudy (loser66):

\[\int_0^1 (9x^3\sqrt{x} -2x^5+e^{-2x}+11x)dx\] is it right?

OpenStudy (anonymous):

yes

OpenStudy (loser66):

so? what stops you work on it?

OpenStudy (loser66):

\(9x^3\sqrt{x} = 9x^3x^{1/2}= 9x^{7/2}\)

OpenStudy (anonymous):

just trying to see other peoples working out. I got (23/3) -(1/2e^2)

OpenStudy (loser66):

correct if e^2 is in denominator \(\dfrac{1}{2e^2}\)

OpenStudy (anonymous):

To help find the velocity of particles requires the evaluation of the indenite integral of the acceleration function, a(t), i.e. v = Z \[\int\limits_{}^{} a(t) dt\] Evaluate the following indenite integrals: (a) \[\int\limits_{?}^{?} \frac{ \ln t }{ t^4 } dt \] (b) \[\int\limits_{?}^{?} (t^2 + 1) \sin9t dt\] sorry you have been really helpful can you please show me the working out for this?